OpenStudy (anonymous):

the formula C=(5/9)(F-32) fives degrees Celsius as a function of degrees Fahrenheit. Find the inverse of this function to convert degrees Celsius to Fahrenheit and use it to find 16 degrees Fahrenheit

OpenStudy (turingtest):

C=(5/9)(F-32) (9/5)C=F-32 (9/5)C+32=F plug in C=16 and you will have your answer

OpenStudy (anonymous):

\[(9/5)16+32 =\frac{304}{5} \]

OpenStudy (turingtest):

I wanted them to do a LITTLE work themselves...

OpenStudy (anonymous):

i used yours